OpenStudy (anonymous):

how do you simplify 2(p+3)-8p+13?

OpenStudy (anonymous):

Apply distributive property to 2(p+3). Do you know how to do that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oour math teacher teaches us nothing

OpenStudy (anonymous):

Distributive property means that a(b+c)=ab+ac

OpenStudy (anonymous):

we did not even learn that

OpenStudy (anonymous):

So 2(p+3) equals 2p+2(3) or 2p+6

OpenStudy (anonymous):

Can you do it from there?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

no we cant we just dont have a clue on simplify

OpenStudy (anonymous):

2p+6-8p+13 It's just adding and subtracting from there on. 2p-8p=-6p 6+13=19 -6p+19 is your answer

OpenStudy (anonymous):

thank you we get that now

OpenStudy (anonymous):

np.